### v3

# An Overview of Prototype Formulations for Interpretable Deep Learning

Abstract
Prototypical part networks offer interpretable alternatives to black-box deep learning models. However, many of these networks rely on Euclidean prototypes, which may limit their flexibility. This work provides a comprehensive overview of various prototype formulations. Experiments conducted on the CUB-200-2011, Stanford Cars, and Oxford Flowers datasets demonstrate the effectiveness and versatility of these different formulations.
XAI, Prototype Learning, Image Classification, Deep Learning
1 Introduction
Deep Learning has achieved high accuracy in many computer vision tasks. However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging. Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of such deep models and increase trust in their decisions.
XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM (Selvaraju et al., 2017) offer explanations for predictions without requiring retraining. While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision-making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019). Second, inherently interpretable methods provide built-in, case-based reasoning processes. For instance, small decision trees are inherently interpretable because their reasoning can be easily understood as a series of if-else statements (Molnar, 2020). However, they are constrained in their representational power.
Figure 1: Different Prototype Formulations. HyperPG is a novel formulation for probabilistic prototypes on a Hypersphere.
Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer. Each neuron in this layer represents a prototype, storing a latent feature vector. The model’s predictions are based on the distances between sample features and prototype parameters, for example, by computing the $L_{2}$-distance. This makes prototype learning essentially a clustering task in latent space (Zhou et al., 2022). Different prototype formulations, such as prototypes using the cosine similarity, induce a hyperspherical structure to the latent space with performance advantages to classification (Mettes et al., 2019). Probabilistic formulations such as Gaussian prototypes enable further downstream tasks relying on probability values such as Bayesian approaches.
With the increased popularity of prototypical part models, there has been a corresponding rise in the number of architectural choices. However, to the best of our knowledge, there is no comprehensive overview of the design choices and prototype formulations for learning interpretable prototypes and their effect on predictive performance.
Table 1: Overview of 14 related models and their prototype learning configurations. 8/14 models use Euclidean prototypes.
Model Similarity Shape Assignment Clf. Head
Prototype Decoder Euclidean Entire Image Class Exclusive FCL Li et al. (2018)
ProtoPNet Euclidean Patch Class Exclusive FCL Chen et al. (2019)
Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL Donnelly et al. (2021)
ProtoPShare Euclidean Patch Merged after Training FCL Rymarczyk et al. (2020)
ProtoPool Euclidean Patch Pooling Shared FCL Rymarczyk et al. (2022)
TesNet Grassman Patch Class Exclusive FCL Wang et al. (2021)
ProtoTree Euclidean Patch Shared Decision Tree Nauta et al. (2021)
ProtoKNN Cosine Patch Class Exclusive KNN Clf Ukai et al. (2023)
PIPNet Cosine Patch Shared FCL Nauta et al. (2023)
LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation Pach et al. (2024)
ProtoSeg Cosine Patch Class Exclusive FCL Zhou et al. (2022)
ProtoGMM Gaussian Patch Class Exclusive FCL Moradinasab et al. (2024)
MGProto Gaussian Patch Class Exclusive Bayesian Likelihood Wang et al. (2024b)
ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration Xue et al. (2024)
This paper explores different prototype formulations. Its contributions are as follows:
Prototype Formulation Overview: A comprehensive overview of different, learnable prototype definitions, including both point based and probabilistic prototypes as shown in Figure 1.
HyperPG: A new probabilistic prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and variance $\sigma^{2}$. This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere. HyperPG’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
Benchmarking Protoypes: Extensive image classification experiments based on prototypical part network architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
2 Related Work
Prototype Learning. In image classification, prototype learning approaches using autoencoders provide high interpretability by reconstructing learned prototypes from latent space back to the image space (Li et al., 2018). However, these approaches are limited in their performance because each prototype must represent the entire image. ProtoPNet (Chen et al., 2019) introduced the idea of prototypical parts. In this setting, each prototype is a latent patch of the input image, commonly a $1\times 1$ latent patch. The prototypes are each associated with a single class and learned via backpropagation without additional information. The similarity of the prototypes to the image patch is based on the Euclidean distance.
Multiple successors build on the idea of ProtoPNet. ProtoPShare (Rymarczyk et al., 2020) merges the class-exclusive prototypes to class-shared ones. ProtoPool (Rymarczyk et al., 2022) directly learns class-shared patch prototypes and pools them by learning a slot assignment. Deformable ProtoPNet (Donnelly et al., 2021) learns a mixture of prototypical parts with dynamic spatial arrangement. For computing the similarity between the prototypes and image patches, Deformable ProtoPNet uses the cosine similarity. TesNet (Wang et al., 2021) proposes to compute the prototype similarity on the Grassman Manifold.
All these approaches share the use of a linear classification layer. However, ProtoTree (Nauta et al., 2021) builds a Decision Tree on the learned Euclidean prototypes and ProtoKNN (Ukai et al., 2023) proposes to use a $k$-nearest-neighbor classifier with cosine prototype similarities.
Recent work proposes to learn more interesting or robust features for the prototypes from the image encoders. PIPNet (Nauta et al., 2023) use augmentations during training to align the learned prototypes with more meaningfull image content. LucidPPN (Pach et al., 2024) propose a hybrid architecture with texture only grayscale prototypes and low resolution color only prototypes.
Zhou et al. (2022) propose to view prototype learning as a clustering task in the network’s latent space for image segmentation. In this setting, the clustering can be done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes. Later work such as ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2024b) build on this idea and model the prototypes as Gaussian distributions.
ProtoPFormer (Xue et al., 2024) adapts Euclidean prototypical part networks to Transformer based architectures by using a VisionTransformer (ViT, Dosovitskiy, 2020) backbone. To avoid the distraction problem due to the global attention mechanism of ViT architectures, ProtoPFormer uses a hybrid approach with global and local prototypes. The local prototypes are encouraged to focus on different parts of the image by modeling the spatial prototype activations as a 2D Gaussian.
Table 1 presents a systematic overview of the discussed works. The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023). There is also a lack for probabilistic prototypes in a hyperspherical space. We introduce HyperPG, which combines the predictive performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
3 Overview: Prototype Formulations
Figure 2: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. Euclidean prototypes compute the $L_{2}$ distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF). HyperPG prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
Prototype Learning is an inherently interpretable machine learning method. The reasoning process is based on the similarity scores of the inputs to the prototypes, retained representations of the training data. For example, a K-Nearest Neighbor (KNN) model is a prototype learning approach with the identity function for representation and an unlimited number of prototypes. In contrast, a Gaussian Mixture Model (GMM) uses a mean representation but restricts the number of prototypes to the number of mixture components.
Prototype learning for deep neural networks involves finding structures in latent space representations. This section provides an overview of existing methods, which are also illustrated in Figure 2. Prior work uses point-based prototypes, computing similarity scores relative to a single point in latent space. On the other hand, probabilistic formulations such as Gaussian prototypes or HyperPG prototypes allow the model to adapt to the variance in the training data.
3.1 Point Based Prototypes
The general formulation of prototypes, as defined in previous work (e.g., Chen et al., 2019), is discussed first. Let $\mathcal{D}=[{\bm{X}},{\bm{Y}}]=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ denote the training set, e.g., a set of labeled images, with classes $C$. Each class $c\in C$ is represented by $Q$ many prototypes ${\bm{P}}_{c}=\{{\bm{p}}_{c,j}\}_{j=1}^{Q}$.
Some feature encoder $\mathrm{Enc}$ projects the inputs into a $D$-dimensional latent space $\mathcal{Z}$, with
${\bm{z}}_{i}=\mathrm{Enc}({\bm{x}}_{i})$ being a feature map of shape $\zeta_{w}\times\zeta_{h}\times D$ with spatial size $\zeta=\zeta_{w}\zeta_{h}$. Commonly, the prototypes ${\bm{p}}$ are also part of $\mathcal{Z}$ with shape $\rho_{w}\times\rho_{h}\times D$, i.e., spatial size $\rho=\rho_{w}\rho_{h}$.
Early approaches based on Autoencoder architectures (Li et al., 2018) use $\rho=\zeta$, meaning the prototype represents the entire image and can therefore be decoded out from latent space back to the input space. Part-based approaches like ProtoPNet and segmentation models like ProtoSeg use $\rho=1$ (Chen et al., 2019; Zhou et al., 2022), meaning each prototype represents some part of the image. Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has is a $\rho=3\times 3$ arrangement of smaller patches, and MCPNet (Wang et al., 2024a), which learns concept prototypes across the latent features.
The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$. For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$. The following equations can be adapted for higher spatial dimensions by summing over the height and width $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each patch of the latent map.
3.1.1 Euclidean Prototypes
ProtoPNet’s Euclidean prototypes leverage the $L_{2}$ similarity. The $L_{2}$ similarity measure is defined as
$s_{L_{2}}({\bm{z}}|{\bm{p}})=\log\left(\frac{\|{\bm{z}}-{\bm{p}}\|^{2}_{2}+1}{\|{\bm{z}}-{\bm{p}}\|^{2}_{2}+\epsilon}\right)$ (1)
and is based on the inverted $L_{2}$ distance between a latent vector ${\bm{z}}$ and a prototype vector ${\bm{p}}$. This similarity measure starts to perform worse with higher numbers of dimensions, as in a large enough space all points are distant to each other and the $L_{2}$ distance looses meaning.
3.1.2 Cosine Prototypes
Prototype formulations based on the cosine similarity create a hyperspherical space. They have been shown to perform well in classification tasks (Mettes et al., 2019) and recent works have started to apply them (e.g., Donnelly et al., 2021; Zhou et al., 2022; Nauta et al., 2021; Ukai et al., 2023). The cosine similarity is defined as
$s_{\cos}({\bm{z}}|{\bm{p}})=\frac{{\bm{z}}^{\top}{\bm{p}}}{\|{\bm{z}}\|_{2}\|{\bm{p}}\|_{2}},$ (2)
which is based on the angle between two normalized vectors of unit length. By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere. The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions. Without the normalization to unit length, we can define a scaled dot-product prototype inspired by the Attention mechanism (Vaswani et al., 2017). However, this similarity measure is not bounded to any interval, as the vector length is proportionally preserved. For more details see Appendix C. Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing two vectors directly.
Both the $L_{2}$ and cosine similarity have been used for classification tasks. The similarity scores are processed by a fully connected layer (e.g., Chen et al., 2019; Donnelly et al., 2021), or a winner-takes-all approach assigns the class of the most similar prototype (e.g., Sacha et al., 2023). The prototypes can be learned by optimizing a task-specific loss, such as cross-entropy, via backpropagation. Alternatively, Zhou et al. (2022) propose “non-learnable” prototypes, whose parameters are obtained via a clustering operation in the latent space instead of backpropagation.
3.2 Probabilistic Prototypes
Viewing prototype learning as clustering problem, point-based prototypes compute the similarity of a latent vector to the center coordinate of each cluster, as represented by the prototype vector. Probabilistic prototypes aim to model the cluster as a probability distribution. Future work could use probabilistic prototypes for extended downstream tasks such as outlier detection or interventions by sampling from generative distributions.
3.2.1 Gaussian Prototypes
Gaussian prototypes model the clusters in latent space as a Gaussian distribution with mean and covariance. By adapting the covariance matrix to the training data a Gaussian prototype with a wide covariance can still have a relatively high response even for larger $L_{2}$ distances from the mean vector. On the other hand, with a very small covariance matrix a Gaussian prototype could show no response unless the mean value is met nearly exactly.
Let the formal definition of a Gaussian prototype be ${\bm{p}}^{G}=(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. The parameters of ${\bm{p}}^{G}_{c,j}$ now track both the mean and covariance of latent vector distribution ${\bm{Z}}_{c,j}$. Each Gaussian prototype ${\bm{p}}^{G}$ thus defines a multivariate Gaussian Distribution ${\mathcal{N}}(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. Gaussian prototypes can be trained using EM for clustering in the latent space (Zhou et al., 2022; Moradinasab et al., 2024; Wang et al., 2024b) or naively by directly optimizing the parameters via Backpropagation. The similarity measure of Gaussian prototypes is defined as the probability density function (PDF) for $D$-dimensional multivariate Gaussians, namely
$s_{\mathrm{Gauss}}({\bm{z}}|{\bm{p}}^{G})$ $=\mathcal{N}({\bm{z}};\bm{{\mu}},\bm{\mathrm{\sigma}})$ (3)
$=\frac{1}{(2\pi)^{\frac{D}{2}}|\bm{\mathrm{\Sigma}}|^{\frac{1}{2}}}$ $\exp\left(-\frac{1}{2}({\bm{z}}-\bm{{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}({\bm{z}}-\bm{{\mu}})\right).$ (4)
It is worth noting that the term inside the exponent is based on the $L_{2}$ distance between the latent vector ${\bm{z}}$ and the mean $\bm{{\mu}}$. This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space. B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space. However, learning a full covariance matrix increases computational requirements, especially with EM clustering.
(a) $\mu=1$
(b) $\mu=0.5$
(c) $\mu=0$
Figure 3: HyperPG prototypes learn a distribution of cosine similarities. They are parameterized by a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and scalar variance $\sigma^{2}$. HyperPG projects a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
Figure 4: Prototype Learning Architecture. The HyperPG module can be easily exchanged to other prototype formulations such as Euclidian or Cosine prototypes. HyperPG uses a Gaussian distribution as Density Estimator, but other PDFs are possible.
3.2.2 Gaussian Prototypes on the Hypersphere - HyperPG
Just as Gaussian prototypes provide a probabilistic formulation for Euclidean prototypes, we aim to develop a probabilistic formulation for hyperspherical prototypes based on the cosine similarity. We name this new formulation Prototypical Gaussians on the Hypersphere (HyperPG). HyperPG prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPG prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPG’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then, HyperPG’s similarity measure based on the truncated Gaussian distribution is defined as
$s_{\mathrm{HyperPG}}({\bm{z}}|{\bm{p}}^{H})$ $=\mathcal{T}_{G}(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma,-1,1)$ (5)
$=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma)}.$ (6)
Figure 3 illustrates the activations of HyperPG’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[0,1]$. The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow. The learned Gaussian distribution of cosine similarities is projected onto the hypersphere’s surface with std $\sigma$ governing the spread of the distribution, and mean $\mu$ the expected distance to the anchor $\bm{{\alpha}}$. For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction. This corresponds to the von Mises-Fisher (vMF) distribution (Appendix B). For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation.
These ring-like activation patterns would require an infinite mixture of prototype vectors pointing in all directions in this hyperplane. HyperPG achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters. This significantly increases HyperPG’s representational power compared to Gaussian prototypes without increasing the computational complexity. This potential capability is also a major difference to the vMF distribution.
HyperPG can be easily adapted to other probability distributions with additional desirable properties. Possible candidate distributions are elaborated on in Appendix B. Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPG idea transferable to other manifolds and applications outside of prototype learning.
4 Training
The original ProtoPNet implementation uses three loss functions: a task specific loss like crossentropy for classification, a cluster loss to increase compactness within a class’s cluster, and a separation loss to increase distances between different prototype clusters.
4.1 Prototype Losses
ProtoPNet defines a cluster loss function to shape the latent space such that all latent vectors ${\bm{z}}_{c}\in{\bm{Z}}_{c}$ with class label $c$ are clustered tightly around the semantically similar prototypes ${\bm{p}}_{c}\in{\bm{P}}_{c}$. The cluster loss function is defined as
$L_{\mathrm{Clst}}=-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|C|}\sum_{c\in C}\max_{{\bm{p}}_{c}\in{\bm{P}}_{c}}\max_{{\bm{z}}_{c,i}\in{\bm{Z}}_{c,i}}s({\bm{p}}_{c},{\bm{z}}_{c,i}),$ (7)
where $s(\cdot,\cdot)$ is some similarity measure. The $L_{\mathrm{Clst}}$-Loss function increases compactness by increasing the similarity between prototypes ${\bm{p}}_{c}$ and latent embeddings ${\bm{z}}_{c}$ of class $c$ over all samples.
An additional separation loss increases the margin between different prototypes. The separation loss function is defined as
$L_{\mathrm{Sep}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|C|}\sum_{c\in C}\max_{{\bm{p}}_{\neg c}\notin{\bm{P}}_{c}}\max_{{\bm{z}}_{c,i}\in{\bm{Z}}_{c,i}}s({\bm{p}}_{\neg c},{\bm{z}}_{c,i}),$ (8)
The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space. Please note, the original ProtoPNet paper uses a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
4.2 Multi-Objective Loss Function
To train a prototype learning network like HyperPGNet for downstream tasks like image classification, a multi-objective loss function is employed. This multi-objective loss function is defined as
$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}},$
where $L_{\mathrm{CE}}$ is the cross-entropy loss over network predictions and ground truth image labels. Our experiments uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ as proposed by ProtoPNet (Chen et al., 2019).
4.3 Network Architecture
For our experiments, we use the network architecture proposed by ProtoPNet (Chen et al., 2019), which we deem the basic deep prototype learning architecture for image classification. Figure 4 illustrates this network architecture with the HyperPG prototype module. In the encoding stage, a pretrained image encoder such as a ResNet model (He et al., 2016) serves as model backbone. Then, a convolution neck projects the high dimensional feature map of the backbone into the lower dimensional prototype space. This neck consists of two $1\times 1$ convolution layers with ReLU activation in between followed by a Sigmoid activation.
In the prototype stage, the prototype module computes the similarity scores. The original ProtoPNet uses Euclidian prototypes, which compute the $L_{2}$ distance between the encoded feature maps and the prototype parameters. Figure 4 shows the make up of the HyperPG module: First, the prototype module computes the cosine similarity of the learnable HyperPG anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck. Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer. Future extensions could adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
The final classification stage produces the output logits based on the prototype module’s similarity scores. ProtoPNet uses a single linear layer due to its inherent interpretability.
4.3.1 Comparison to Recent Works
Our experiments use only the basic ProtoPNet architecture and focus on the effect of different prototype formulations. This section provides a short overview of the required changes to this basic architecture in order to implement the models described in Table 1. LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder specializing in color-only and texture-only feature extraction. Deformable ProtoPNet (Donnelly et al., 2021) makes changes towards the input of the prototype module to allow for dynamic spatial arrangements of the prototypes, i.e., “deformations”. ProtoPool (Rymarczyk et al., 2022) makes changes towards the output of the prototype module, by also learning how to dynamically pool the different prototype activations. ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) propose to use a different inherently interpretable model for the classification stage. ProtoPFormer (Xue et al., 2024) builds exclusively on a Vision Transformer backbone and uses different network branches for the global information encoded in the class token and the local information in the image tokens.
5 Experiments
The experiments were performed on CUB-200-2011 (CUB) with 200 bird species (Wah et al., 2011) with ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) backbones. Later experiments also explore the Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets. Implementation details such as data preprocessing or model implementation are provided in Appendix D.
Table 2: Mean Top-1 Test Accuracies and Std for different prototype formulations with ResNet50 Backbones averaged over 3 random seeds.
ResNet50 Dense121 ViT-B16
Baseline 79.2 $\pm$ 0.2 75.5 $\pm$ 0.2 78.5 $\pm$ 0.8
Euclidian (ProtoPNet) 35.3 $\pm$ 1.4 33.6 $\pm$ 3.0 9.4 $\pm$ 0.4
Gaussian 61.1 $\pm$ 0.2 45.9 $\pm$ 5.1 4.2 $\pm$ 1.3
Cosine 71.7 $\pm$ 0.4 68.7 $\pm$ 0.5 69.2 $\pm$ 3.9
HyperPG (Ours) 74.3 $\pm$ 0.6 70.7 $\pm$ 0.3 72.9 $\pm$ 0.1
5.1 Quantitative Results
Table 2 reports the mean top-1 accuracy and standard deviation for the tested models on the CUB dataset with backbone models ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) pretrained on ImageNet. Please note that in contrast to prior work, we do not report scores from model ensembles or using pretrained iNaturalist weights. The black-box baselines perform the best across all experiments, which we see as a theoretical upper bound for the performance. The classical Euclidean prototypes perform the worst across all experiments. A first model improvement can be achieved by switching to probabilistic Gaussian prototypes. The hyperspherical Cosine prototypes further improve the performance. HyperPG combines the advantages of Gaussian and Cosine prototypes, further closing the gap to the upper bound presented by the black-box baselines. Notably, both Euclidian and Gaussian prototypes are unable to learn on the latent features of ViT-B16. We theorize that the attention mechanism of the Transformer architecture produces a near-hyperspherical latent space. This highlights the importance of choosing the right prototype formulation for the used backbone. For example, language aligned models such as (Radford et al., 2021) use the cosine similarity for their alignment. We expect that only hyperspherical prototypes can produce meaningful results in this setting.
Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone. Mean and std over 3 random seeds.
Figure 5 shows the test accuracy over epochs for CUB with ResNet50 backbone. The standard Euclidean prototypes take the longest time (200 Epochs) until convergence while also achieving the lowest accuracy. By changing only the prototype formulation to a Gaussian one, the model achieves higher performance and converges at 100 epochs. Both the cosine and HyperPG formulation outperform the other formulations in model accuracy and training time, requiring only 30 training epochs.
5.2 Ablation: Number of Prototypes
Figure 6: Ablation: Test Accuracy on CUB with different numbers of prototypes per class.
We conduct an ablation study over the number of prototypes per class and present the test accuracy scores in Figure 6. Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added. The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
5.3 Ablation: Number of Dimensions
We conduct an ablation study over the number of prototype dimensions and present the test accuracy scores in Figure 7. Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality. The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality. The performance of Euclidean prototypes suffers most with changes in dimensionality.
Figure 7: Ablation: Test Accuracy on CUB with different numbers of prototype dimensions.
5.4 Ablation: Additional Datasets
We perform additional experiments on two different datasets. Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 different species (Nilsback & Zisserman, 2008). The relative differences in performance seem to hold across the different datasets, with Euclidean prototypes performing the worst. Interestingly, Gaussian prototypes are able to match the performance of the hyperspherical formulations more closely.
Table 3: Top-1 Test Accuracy with Resnet50 Backbone.
Cars Flowers
Baseline 85.2 90.0
Euclidean (ProtoPNet) 36.2 15.0
Gaussian 73.0 74.8
Cosine 79.6 85.5
HyperPG (Ours) 79.0 87.8
6 Interpretability Analysis
We perform a qualitative analysis of the different prototype formulations with ResNet50 backbone (see Figure 8). One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM). However, this visualization technique has several shortcomings, which we discuss in more detail in Appendix A. In order to offer an alternative to the PAM visualization, we also present GradCAM (Selvaraju et al., 2017) overlays.
Both Euclidian and HyperPG prototypes do not actually learn class-specific prototypes, in contrast to Gaussian or Cosine prototypes. Interestingly, all prototype formulations seem to activate in similar regions in the test image, but differ with the known training examples. HyperPG could already leverage class-shared prototypes, which would explain its robustness to changes in the number of prototypes per class.
There seems to be no qualitative difference between the visualizations produced with the various prototype formulations using the CNN based backbones. As expected, the ViT backbone (see Figure 9) suffers from the “distraction problem” as observed by ProtoPFormer (Xue et al., 2024). We present additional examples in Appendix A.
Refer to caption Refer to caption
(a) Euclidean (R50)
Refer to caption Refer to caption
(b) Gaussian (R50)
Refer to caption Refer to caption
(c) Cosine (R50)
Refer to caption Refer to caption
(d) HyperPG (R50)
Figure 8: Left: Test Image. Right: 3 closest Training Patches to the Prototype. Top Row: PAM visualization. Bottom Row: GradCAM visualization. R50: Resnet50 Backbone. ViT: ViT-B16 Backbone.
Figure 9: Cosine Protoypes with ViT. The overlayed heatmap show the distraction problem caused by the attention mechanism.
7 Conclusion
This work provides comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in an Euclidean or hyperspherical latent space. We introduce HyperPG, a new prototype representation based on Gaussian distributions on the surface of a hypersphere. Our experiments showcase the differences in performance and robustness to hyperparameters across different datasets and backbone architectures for these prototype formulations. Our results demonstrate the performance advantage of hypersherical prototypes and probabilistic prototypes in contrast to the commonly used Euclidean formulation, with no loss to the inherent interpretability.
Future work could explore further refinements to the probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b) to further enhance interpretability and performance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Extended Interpretability Analysis
A.1 Potential Problems with Prototype Activation Maps
One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM). In this manner, the latent similarity map for each prototype is scaled back up to the original input resolution and overlayed over the image. A box is drawn around the 95th quantile to mark the highest activated image region. This technique has some obvious flaws: The latent similarity map is usually very low resolution, e.g., for ResNet50 $7\times 7$, unless a segmentation model such as DeeplabV3 (Chen et al., 2017) is used. Any perceived gradient in the overlay is likely an artifact from the upscaling operation. Additionally, there exist robustness problems with this visualization (Sacha et al., 2024). By modifying parts of the background or adding noise to the image, the activation map can change, most likely due to the spatial pooling in the convolution pyramids.
Appendix B Adapting HyperPG to other Probability Distributions
We define HyperPG prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and variance $\sigma^{2}$ of cosine similarities around an anchor vector $\alpha$. This idea of learning a distribution of cosine similarity values around an anchor $\bm{{\alpha}}$ can be adapted to other distributions. This sections introduces some potential candidates. As early experiments on the CUB-200-2011 dataset showed no significant difference in performance, these sections are relegated to the appendix.
B.1 Cauchy Distribution
One theoretical disadvantage of the Gaussian distribution is the fast approach to zero, which is why a distribution with heavier tails such as the Cauchy distribution might be desirable. The Cauchy distribution’s PDF is defined as
$\mathcal{C}(x;x_{0},\gamma)=\frac{1}{\pi\gamma\left(1+\left(\frac{x-x_{0}}{\gamma}\right)^{2}\right)},$ (9)
with median $x_{0}$ and average absolute deviation $\gamma$. The HyperPG prototypes with Cauchy are defined as accordingly as ${\bm{p}}^{\mathrm{Cauchy}}=(\bm{{\alpha}},x_{0},\gamma)$.
Figure 18 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$. The Gaussian distributions PDF quickly approaches zero and stays near constant. This could potentially cause vanishing gradient issues during training. The heavier tails of the Cauchy distribution ensure that for virtually the entire value range of the cosine similarity, gradients could be propagated back through the model. However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPG with the Gaussian or Cauchy distribution.
Figure 18: PDF for the Gaussian and Cauchy distribution of cosine similarity values. The Cauchy distribution has heavier tails, avoiding vanishing gradients issues.
B.2 Truncated Distributions
The cosine similarity is defined only on the interval $[-1,1]$. This makes it attractive to also use truncated probability distributions, which are also only defined on this interval. The truncation imposes a limit on the range of the PDF, thereby limiting the influence of large values for the distribution’s $\sigma$ or $\gamma$ parameter, respectively. The truncated Gaussian pdf $\mathcal{T}_{\mathrm{Gauss}}$ requires the cumulative probability function $\mathcal{G}$ and error function $f_{\mathrm{err}}$, and is defined as
[…]
${\mathcal{N}}(x;\mu,\sigma)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right),$ (11)
$\mathcal{G}(x,\mu,\sigma)$ $=\frac{1}{2}\left(1+f_{\mathrm{err}}\left(\frac{x-\mu}{\sigma\sqrt{2}}\right)\right),$ (12)
$\mathcal{T}_{\mathrm{Gauss}}(x,\mu,\sigma,a,b)$ $=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma).},$ (13)
with lower bound $a$ and upper bound $b$, e.g., for the cosine similarity $a=-1$ and $b=1$. Similarly, the truncated Cauchy distribution can be applied, which is defined as
$\mathcal{T}_{\mathrm{Cauchy}}(x,x_{0},\gamma,a,b)=\frac{1}{\gamma}\left(1+\left(\frac{x-x_{0}}{\sigma}\right)^{2}\right)^{-1}\left(\arctan\left(\frac{b-x_{0}}{\gamma}\right)-\arctan\left(\frac{a-x_{0}}{\gamma}\right)\right)^{-1}.$ (14)
Figure 19 shows the test accuracy of three HyperPGNet models with Gaussian, truncated Gaussian and truncated Cauchy distribution on the CUB-200-2011 dataset. While difference in test performance and learning speed were minimal on the CUB-200-2011 dataset, further exploration is necessary, as other experiments showed that the concept-alignment on CUB-200-2011 dominates the learning process, lessening the influence of the prototypes.
Figure 19: CUB-200-2011 Test Accuracy with HyperPG prototypes using different PDFs.
B.3 von Mises-Fisher Distribution
The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere (Hillen et al., 2017). The density function $f_{d}$ of the vMF distribution for a $D$-dimensional unit-length vector ${\bm{v}}$ is defined as
$f_{d}({\bm{v}}|\bm{{\alpha}},\kappa)=C_{d}(\kappa)\exp\left(\kappa\bm{{\alpha}}^{\top}{\bm{v}}\right),$ (15)
with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$. The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2025-01-10 only supports the vMF distribution for $D\leq 5$.
However, the vMF distribution is a viable similarity measure when using the unnormalized density function with $C_{d}(\kappa)=1$. Working with unnormalized densities highlights the relationship between the normal distribution and the vMF distribution.
Let $\hat{G}$ be the unnormalized PDF of a multivariate Gaussian with normalized mean $\bm{{\alpha}}$ and isotropic covariance $\bm{{\sigma}}^{2}=\kappa^{-1}{\bm{I}}$, then it is proportional to the vMF distribution for normalized vectors ${\bm{v}}$ with $|{\bm{v}}|=1$, as shown by
$\hat{G}({\bm{v}}|\bm{{\alpha}},\kappa)$ $=\exp\left(-\kappa\frac{({\bm{v}}-\bm{{\alpha}})^{\top}({\bm{v}}-\bm{{\alpha}})}{2}\right)$ (16)
$=\exp\left(-\kappa\frac{{\bm{v}}^{\top}{\bm{v}}+\bm{{\alpha}}^{\top}\bm{{\alpha}}-2{\bm{v}}^{\top}\bm{{\alpha}}}{2}\right)$ (17)
$=\exp\left(-\kappa\frac{1+1-2{\bm{v}}^{\top}\bm{{\alpha}}}{2}\right)$ (18)
$=\exp\left(-\kappa\frac{2-2{\bm{v}}^{\top}\bm{{\alpha}}}{2}\right)$ (19)
$=\exp\left(-\kappa\frac{1-{\bm{v}}^{\top}\bm{{\alpha}}}{1}\right)$ (20)
$=\exp\left(\kappa({\bm{v}}^{\top}\bm{{\alpha}}-1)\right)$ (21)
$=\exp(\kappa{\bm{v}}^{\top}\bm{{\alpha}}-\kappa)$ (22)
$=\exp(\kappa)^{-1}\exp(\kappa{\bm{v}}^{\top}\bm{{\alpha}})$ (23)
$\sim\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (24)
Equation 21 also shows the relationship to the HyperPG similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$. Figure 20 presents a simulation of the vMF distribution on a 3D sphere. While both the vMF distribution and HyperPG activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Figure 3. The ring pattern produced by adapting HyperPG’s mean similarity $\mu$ could be approximated by a mixture of vMF distributions.
Figure 20: Changing the concentration parameter $\kappa$ is akin to changing HyperPG’s std $\sigma$.
B.4 Fisher-Bingham Distribution
As the vMF distribution is the equivalent of an isotropic Gaussian distribution on the surface of a hypersphere, the Fisher-Bingham (FB) distribution is the equivalent of a Gaussian with full covariance matrix. Similar to the vMF, the normalization constant is difficult to compute for higher dimensions, but the unnormalized density function remains feasible.
For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$. The FB’s unnormalized PDF is defined as
$b({\bm{v}}|{\bm{A}},\kappa,\beta)=\exp\left(\kappa\bm{{\alpha}}_{1}^{\top}{\bm{v}}+\sum_{j=2}^{D}\beta_{j}\left(\bm{{\alpha}}_{j}^{\top}{\bm{v}}\right)^{2}\right).$ (25)
The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Figure 21. However, the parameter count and constraints are higher.
Figure 21: Illustration of the Fisher-Bingham Distribution in $D=3$
B.5 Mixture Models
HyperPG’s probabilistic nature lends itself to a mixture formulation. Let the definition of a HyperPG Mixture Prototype be ${\bm{p}}^{M}=(\bm{{\alpha}},\mu,\sigma,\pi)$ with additionally learned mixture weight $\pi$. Further, let’s define the probability of a latent vector ${\bm{z}}$ belonging to a Gaussian HyperPG prototype ${\bm{p}}$ as
$\phi({\bm{z}}|{\bm{p}})=s_{\mathrm{HyperPG}}({\bm{z}}|{\bm{p}}).$ (26)
Then the probability of ${\bm{z}}$ belonging to class $c$ can be expressed through the mixture of all prototypes ${\bm{p}}_{c}\in{\bm{P}}_{c}$ of class $c$, i.e.,
$\phi({\bm{z}}|c)=\sum_{{\bm{p}}_{c}\in{\bm{P}}_{c}}\pi({\bm{p}}_{c})\phi({\bm{z}}|{\bm{p}}_{c}).$ (27)
First experiments with mixture of HyperPG prototypes did not show any improvement over the standard formulation. However, this might change with other datasets.
Appendix C Scaled Dot-Product Prototype
This section does not reiterate the full attention mechanism. We recommend to the interested reader to refer to the original work (Vaswani et al., 2017). Rather, we will take a high-level view of the main equation and explain how we apply it to a prototype formulation.
Vaswani et al. (2017) define the scaled dot-product attention for some input matrices query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ as
$\mathrm{Attention}({\bm{Q}},{\bm{K}},{\bm{V}})=\mathrm{softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{k}}}\right){\bm{V}}$ (28)
with $d_{k}$ being the number of dimensions for the key matrix ${\bm{K}}$.
We can use the scaled dot-product to define a new prototype similarity function
$s_{\mathrm{sdot}}({\bm{z}}|{\bm{p}})=\frac{{\bm{z}}^{\top}{\bm{p}}}{\sqrt{D}}$ (29)
for a latent patch ${\bm{z}}$ and prototype vector ${\bm{p}}$ with dimensionality $D$. In contrast to the cosine similarity, the vectors are not normalized to unit length, which preserves more information. However, dividing by the square root of the number of dimensions $D$ acts as a heuristic regarding vector length. Intuitively we can think about it as instead of a hard projection onto the surface of the unit-radius hypersphere, all points are moved towards the surface. Figure 22 illustrates this scaling. In this manner, the angle between vectors becomes a meaningful measurement, while also retaining some information encoded by the vector length.
Figure 22: Illustration of Cosine Similarity and Scaled Dot-Product
Appendix D Implementation Details
D.1 Data Preprocessing
In contrast to prior work (e.g. Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch. The input images were cropped to the bounding box annotations, then resized to a resolution of $224\times 224$. The augmentations consisted of RandomPerspective, RandomHorizontalFlip and ColorJitter.
D.2 Hyperparameters
The prototypical part networks ProtoPNet and HyperPGNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 128$. The models were trained with a minibatch size of 48 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
D.3 Compute Resources
All experiments were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM) per model. On CUB-200-2011 the training duration was roughly 40 seconds per epoch, regardless of the used prototype formulation.
D.4 Model Implementation
The experiments used various feature encoding backbones such as ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) with pretrained weights on ImageNet. All models use a single linear output layer as a classification head.
The prototypical networks follow the architecture proposed by ProtoPNet (Chen et al., 2019): After the pretrained feature encoder, a bottleneck consisting of a Convolution Layer, ReLU activation, Convolution Layer and Sigmoid activation reduces the dimensions of the feature map to the prototype space. For example, the ResNet50 encoder produces a feature map of $H\times W\times D$ shape $7\times 7\times 2048$. The bottleneck produces a dimension reduction to shape $7\times 7\times 128$.
The prototype modules produce the prototype similarity scores, which are passed through a global max pooling, before being passed to the classification head.
Appendix E Parameter Distribution of HyperPG
To evaluate, if HyperPG actually uses the probabilstic components, we look at the learned mean $\mu$ and standard deviation $\sigma$ values after training on CUB200-2011.
Figure 23 plots the mean and sigma values. The red star marks the theoretical solution if it would learn something similar to the pure Cosine prototypes. The pattern demonstrates, that HyperPG actually learns different parameters for the Gaussian components of each prototype. Interestingly, two modes at $\sigma=0.5$ and $\sigma=0.4$ appear with a distinct line shape, which warrant future exploration.
Figure 23: HyperPG learned mean and sigma values.